OpenStudy (konradzuse):

Find the extrema and the points of inflection (if any exist) of the function. Use a graphing utility to graph the function and confirm your results. (Round your answers to three decimal places. Order your answers from smallest to largest x, then from smallest to largest y.) g(x) = 1/√2π^(e−(x − 5)2/2)

OpenStudy (konradzuse):

IT looks like that.

OpenStudy (konradzuse):

2pi is under the sqrt.

OpenStudy (konradzuse):

it's a weird looking problem lol.

OpenStudy (turingtest):

\[\huge \frac1{\sqrt{2\pi}^{\left(e-{(x-5)^2\over2}\right)}}\]

OpenStudy (konradzuse):

yeah...

OpenStudy (konradzuse):

I don't like it at all!

OpenStudy (konradzuse):

according to the practice version it seems I take the derivative twice to find the data, it's soo konfusing... All of these questions are weird... :(.

OpenStudy (turingtest):

ok this is very ugly so I want you to be \(sure\) that the problem is\[\huge \frac1{\sqrt{2\pi}^{\left(e-{(x-5)^2\over2}\right)}}\]everything is under 1 and all that is correct?

OpenStudy (konradzuse):

sec adding an attachment.

OpenStudy (konradzuse):

ugly....

OpenStudy (turingtest):

oh that's just terribly written... what the heck is that 2/2 ??? is the exponent 2/2=1 or do they means that whole part is over 2? somebody messed up I think in typing this out

OpenStudy (konradzuse):

yeah I was going to say it looks like 2/2 but that would be 1 so ... :(

OpenStudy (turingtest):

\[\huge g(x)=\frac1{\sqrt{2\pi}}e^{\frac{-(x-5)^2}2}\]is what it looks like to me I guess...

OpenStudy (turingtest):

so you wanna take a stab at that derivative? that's how we start this whole thing....

OpenStudy (anonymous):

\[\huge \frac1{\sqrt{2\pi}}{e^{-{x-5^2\over2}}}\] is my guess

OpenStudy (konradzuse):

I'm scared :(.

OpenStudy (konradzuse):

there's like an entire train of chain rules :P.

OpenStudy (anonymous):

oops i should have looked at the picture

OpenStudy (turingtest):

yeah it's chain rule madness

OpenStudy (konradzuse):

Lets see... I would start off with the e^-(x-5)^2 would be 2(x-5)e^(x-5)^2

OpenStudy (turingtest):

unless satellite has some more clever way :D

OpenStudy (anonymous):

not so bad at all. ignore the \(\frac{1}{\sqrt{2\pi}}\) as that s a constant and just take the derivative of \(\frac{-(x-5)^2}{2}\)

OpenStudy (konradzuse):

but there's a chain rule it itself so then there would be a 1 I guess so I guess that doesn't even matter....

OpenStudy (turingtest):

yeah true, not really that bad...

OpenStudy (konradzuse):

then I guess we would need the quotient rule because of that 2?

OpenStudy (anonymous):

the derivative if \(-x+5\) or \(5-x\)

OpenStudy (konradzuse):

or prob not...

OpenStudy (turingtest):

no, quotient rule would be if you had a funciton of x in the denominator

OpenStudy (anonymous):

*is

OpenStudy (konradzuse):

oh okay.

OpenStudy (anonymous):

the two in the denominator is a constant. leave it

OpenStudy (konradzuse):

(2(x-5)e^(x-5)^2)/2?

OpenStudy (konradzuse):

err

OpenStudy (konradzuse):

(-2(x-5)e^-(x-5)^2)/2 ??

OpenStudy (anonymous):

\[\frac{d}{dx}[\frac{-(x-5)^2}{2}=-x+5\] by the power rule

OpenStudy (anonymous):

the twos cancel

OpenStudy (konradzuse):

ah gotcha.

OpenStudy (konradzuse):

I hate this question... :(

OpenStudy (konradzuse):

so then it;s (-x+5)e^-(x+5)^2?

OpenStudy (anonymous):

well there is still a two in the denominator of the exponent

OpenStudy (konradzuse):

ah okay. So that orignal 2 came down as well and then they cancelled gotcha.

OpenStudy (anonymous):

but before you start hating this question, lets do it in our heads

OpenStudy (konradzuse):

I'll hate it forever and ever :(. ;P

OpenStudy (anonymous):

you have the derivative as some constant times \(-x+5\) times e to the power of something, and you want the critical points. e to the power of whatever is never zero, so your only job is to set \(-x+5=0\) and solve, which i assert is not that hard!

OpenStudy (konradzuse):

so it should be 1/sqrt(2pi) -(x+5)e^-(x+5)^2 / 2

OpenStudy (konradzuse):

x = 5?

OpenStudy (konradzuse):

stil have no idea :(

OpenStudy (konradzuse):

and then just solve for y?

OpenStudy (anonymous):

yes the critical point is at \(x=5\)

OpenStudy (konradzuse):

welcome back satellite :P

OpenStudy (konradzuse):

now to solve for y.

OpenStudy (konradzuse):

y = that big function right?

OpenStudy (asnaseer):

where are you stuck on this?

OpenStudy (konradzuse):

how to find y? is it = to the big nasty eq?

OpenStudy (konradzuse):

just plug in for x and just go huh?

OpenStudy (konradzuse):

and I bet I gotta use logs too...

OpenStudy (asnaseer):

yes, the maximum will be found by pluggin x=5 into:\[\huge g(x)=\frac1{\sqrt{2\pi}}e^{\frac{-(x-5)^2}2}\]

OpenStudy (asnaseer):

no logs - note that you will end up with \(e^0\) which is just equal to one

OpenStudy (asnaseer):

do you understand?

OpenStudy (konradzuse):

wouldn't it be 1/2?

OpenStudy (konradzuse):

woops nvm.

OpenStudy (konradzuse):

so we then have 1/(sqrt(2pi) = y?

OpenStudy (asnaseer):

\[g(5)=\frac{1}{\sqrt{2\pi}}\]

OpenStudy (asnaseer):

this represents the maximum value for the function g(x)

OpenStudy (asnaseer):

to find the points of inflection you need to differentiate g(x) twice, equate it to zero, and then solve to find which value(s) of x make it zero

OpenStudy (konradzuse):

g(x) = 1/sqrt(2pi) -(x+5)e^-(x+5)^2 / 2 ??

OpenStudy (asnaseer):

???

OpenStudy (konradzuse):

g(x)'

OpenStudy (konradzuse):

haha :)

OpenStudy (konradzuse):

lemme retype that :P.

OpenStudy (konradzuse):

\[\frac{1}{\sqrt{2\pi}} -(x+5)e^\frac{{-(x+5)^2}}{2}\]

OpenStudy (asnaseer):

you should get:\[g'(x)=\frac{(x-5)e^{-\frac{(x-5)^2}{2}}}{\sqrt{2\pi}}\]

OpenStudy (asnaseer):

now differentiate this again to get the second derivative

OpenStudy (konradzuse):

ok

OpenStudy (konradzuse):

\[\frac{2(x-5)(x-5)e^-\frac{(x-5)^2}{2}}{\sqrt{2\pi}}\] ?

OpenStudy (konradzuse):

But I'm guessing the 2 in front cancels with the 2 like from before...

OpenStudy (asnaseer):

it is actually:\[g''(x)=\frac{(x-5)^2e^{-\frac{(x-5)^2}{2}}}{\sqrt{2\pi}}-\frac{e^{-\frac{(x-5)^2}{2}}}{\sqrt{2\pi}}\]\[\qquad=\frac{((x-5)^2-1)e^{-\frac{(x-5)^2}{2}}}{\sqrt{2\pi}}\]\[\qquad=\frac{(x^2-10x+25-1)e^{-\frac{(x-5)^2}{2}}}{\sqrt{2\pi}}\]\[\qquad=\frac{(x^2-10x+24)e^{-\frac{(x-5)^2}{2}}}{\sqrt{2\pi}}\]\[\qquad=\frac{(x-4)(x-6)e^{-\frac{(x-5)^2}{2}}}{\sqrt{2\pi}}\]

OpenStudy (konradzuse):

What the poop....

OpenStudy (asnaseer):

now find what values of x make this zero

OpenStudy (konradzuse):

so where do we square it, and get the other e business?

OpenStudy (asnaseer):

you need to refresh yourself on how to differentiate. this is using the chain rule.

OpenStudy (konradzuse):

yeah I figured, just trying to figure out from what :P.

OpenStudy (konradzuse):

Well I guess we bring another (x-5) down so that's where the ^2 is? And then we leave the other e^ sfdjdlskfjdslk which causes a - that?

OpenStudy (asnaseer):

roughly - yes

OpenStudy (asnaseer):

here, I entered this into wolframalpha for you: http://www.wolframalpha.com/input/?i=d^2%2Fdx^2+%281%2Fsqrt%282*pi%29%29*e^%28-%28x-5%29^2%2F2%29 click "show steps" to see how the derivative is calculated

OpenStudy (konradzuse):

okies.

OpenStudy (asnaseer):

what is left in this problem is as follows: 1. find the value(s) of x that make g''(x)=0 2. find the corresponding value(s) of g(x) for each solution in step 1 this will give you your point(s) of inflection.

OpenStudy (konradzuse):

that damn product rule! :P

OpenStudy (konradzuse):

would 0 work?

OpenStudy (konradzuse):

and 12?

OpenStudy (konradzuse):

144- (10*12[120]) + 24?

OpenStudy (konradzuse):

or maybe it's 4 and 6 :P

OpenStudy (konradzuse):

yeah it's 4 and 6, and the y's for both are 0.242 :)

OpenStudy (asnaseer):

that is the correct answer - x=4 and x=6 are the x-coordinates of the inflection points. I haven't worked out the y-coordinates but assume you have done it correctly by working out g(4) and g(6).

OpenStudy (asnaseer):

\((x-5)^2\) would give 1 for both x=4 and x=6, so both y values are indeed the same.